OpenStudy (anonymous):

Lim x → 3 (x^3 − 3x + 8) = 26 ε = 0.2 δ = ε = 0.1 δ = I have no idea how to do this sort of problem. Can someone show me how to solve for at least the first delta. I really need to learn it! You would SERIOUSLY be a life safe. plus i'll give a medal! :)

OpenStudy (sirm3d):

\[\|(x^3 - 3x+8)-26|<\varepsilon\\|x^3-3x-18|<\varepsilon\\|(x-3)(x^2+3x+6)|<\varepsilon\]

OpenStudy (sirm3d):

suppose \[0<|x-3|<\delta\leq 1\\-1<x-3<1\\2<x<4\\2^2<x^2<4^2\text{ and }3(2)<3x<3(4)\\4<x^2<16\text{ and } 6<3x<12\\4+6<x^2+3x<16+12\\10+6<x^2+3x+6<28+6\\16<x^2+3x+6<34\\|x^2+3x+6|<34\]

OpenStudy (sirm3d):

now\[|x-3|<\delta,\quad \|x^2+3x+6|<34,\qquad \delta \leq 1\\|x-3|\cdot|x^2+3x+6|<\delta(34),\qquad \delta\leq 1\] put \[34\delta = \varepsilon\\\delta = \text{ minimum}\left(\frac{\varepsilon}{34},1\right)\] note... if \(\varepsilon>34\), choose \(\delta = 1\) otherwise choose \(\displaystyle \delta=\frac{\varepsilon}{34}\)

OpenStudy (sirm3d):

just sub your \(\varepsilon\) value into the equation, and you have your answers.

OpenStudy (anonymous):

Hello, Thanks for taking the time to reply to my question! The answer for this problem ( if epsilon is 0.2) is 0.0083, however I don't know how the answer came to this. 0.1 / 34 isn't the same number?

OpenStudy (sirm3d):

the answer differs from the assumed initial value of \(\delta\), which in my work, i chose \(\delta \leq 1\) if your teacher has suggested a starting value for \(\delta\) then you should arrive at the expected result.

OpenStudy (anonymous):

I understand. The work just says to find the largest possible values of δ that correspond to ε = 0.2 and ε = 0.1

OpenStudy (anonymous):

So you found a distance of 1 between x and the target point. How would we know what the maximum would be?

OpenStudy (sirm3d):

the maximum would be \(\displaystyle \frac{\varepsilon}{34}\)

OpenStudy (anonymous):

that's not the minimum? I'm still confused how the work produced 0.0083 and yet it didn't show a starting point? :(

OpenStudy (sirm3d):

there is no minimum. the solution is given by \[\delta\leq\frac{\varepsilon}{34},\qquad \color{red}{\delta \leq 1}\] the inequality in red produced the 34 in the denominator. a different delta, say \(\color{blue}{\delta\leq 0.1}\) will yield denominator different from 34.

OpenStudy (anonymous):

You've been a great deal of help sir! I've learned a lot here. However, I don't know why the homework would be so vague, just poor direction? :) But as you suggested, using delta < 0.1 gave 0.008 which is close.